OpenStudy (firejay5):

Show work & explain steps! :D Medal will be given! 8.

OpenStudy (firejay5):

\[\frac{ 4 }{ x - 1 } = \frac{ x + 1 }{ 12 }\]

OpenStudy (shamim):

ok

OpenStudy (anonymous):

nvm

OpenStudy (shamim):

i wanna try it

OpenStudy (firejay5):

Do you want to try?

OpenStudy (shamim):

\[\frac{ 4 }{ x-1 }=\frac{ x+1 }{ 12 }\]

OpenStudy (shamim):

or\[(x+1)(x-1)=4 \times 12\]

OpenStudy (shamim):

or\[x ^{2}-1=48\]

OpenStudy (shamim):

or\[x ^{2}=48+1\]

OpenStudy (shamim):

or\[x ^{2}=49\]

OpenStudy (shamim):

or\[x=7, -7\]

OpenStudy (firejay5):

did you take the square root or what

OpenStudy (firejay5):

@ajprincess Is he right?

OpenStudy (ajprincess):

ya he is:) taking square root both sides leave u with that answer:)

OpenStudy (firejay5):

thanks so much @ajprincess I might need you again this afternoon of some more math problems

OpenStudy (shamim):

ya i took square root

OpenStudy (ajprincess):

welcome:) will try my best to help:)

OpenStudy (firejay5):

awesome

OpenStudy (shamim):

i also wanna try to so;veur problem

OpenStudy (shamim):

if i can